# Reward-Augmented Data Enhances Direct Preference Alignment of LLMs

### v3

@@ 6.2 Ablation Studies
We provide two pieces of evidence that
@@
our method can get more juice out of the preference data compared to directly applying DPO. Firstly, we evaluate SPPO (Wu et al., 2024) fine-tuned with DPO on UltraFeedback (UF). The results are shown in Table 4. Since the SPPO model is already trained on UltraFeedback from Gemma-2-9B-It, an additional round of DPO training with the same data significantly degrades its performance. In contrast, performing DPO on Reward-Augmented (RA) UltraFeedback results in substantial performance gains for SPPO, indicating that our method enhances the utility of the preference data.
The second evidence is that after DPO, the implicit reward can be used to relabel and augment the same preference data. Specifically, after training Qwen2-7B-Instruct with DPO on UltraFeedback, we leverage the resulting model $\pi_{\text{DPO}}$ to calculate the implicit reward for each prompt $x$ and response $y$, i.e., $\widehat{r}=\beta(\log\pi_{\text{DPO}}(y\mid x)-\log\pi_{\text{Qwen}}(y\mid x))$. Then we perform DPO on Qwen2-7B-Instruct
@@
using the Implicit-Reward-Augmented (IRA) UltraFeedback. The results are shown in Table 5. We observe that augmenting the data with the implicit reward from the DPO (UF) model leads to superior performance even compared to augmenting the data with reward scores from the LLM judge, i.e., DPO (RA). This result highlights that DPO does not fully exploit the potential of the data. Moreover, this ablation demonstrates that our method is compatible with binarized preference datasets that only contain chosen and rejected response pairs, bypassing the need of judge models.
@@
In this part, we
@@
show that the success of our method is not merely due to the increased size of the training dataset. To illustrate this, we perform DPO on the dataset where reward augmentation is applied to the first half of the UltraFeedback data, which we denote as DPO (Half RA). By doing so, the reward-augmented data is of the same size as the original dataset, but with only half of the prompts and the corresponding responses being utilized. It can be observed from Table 6 that DPO (Half RA) outperforms fine-tuning on the whole UltraFeedback (UF) by a large margin and achieves comparable performance to applying reward augmentation across the entire UF dataset, which is denoted as DPO (RA).
@@
Reward-Augmented Data Mitigates the Unlearning Issue.
We first demonstrate that DPO suffers from the limitation of unnecessarily unlearning high-quality rejected responses, as discussed in Section 3.1. Specifically, on the test set of UltraFeedback, we calculate the log probability of each rejected response for the Qwen2-7B-Instruct model, its DPO (UF) model, and our method DPO (RA). In Figure 3, we plot the expected log probability for rejected responses with reward scores $\geq 5$. We find that DPO substantially decreases the probability of these high-quality rejected responses, confirming that the unlearning issue arises in practice. In contrast, our method alleviates this issue, although the probability is still slightly lower than the base model, which is proven to be DPO’s feature (Rafailov et al., 2024a; Zhang et al., 2024; Xu et al., 2024b).
Impact of the Accuracy of AI Feedback.
Llama-3- PairRM (0.4B) ArmoRM (8B)
8B-Instruct
DPO (UF)
DPO (RA+)
DPO (RA)
DPO (UF)
DPO (RA+)
LC WR 22.92 41.76 44.72 48.20 42.32 48.73
WR 23.15 45.79 44.70 53.17 42.79 45.36
SLiC-HF ORPO CPO RRHF KTO IPO RPO R-DPO SimPO Ours
LC WR 26.9 28.5 28.9 31.3 33.1 35.6 40.8 41.1 44.7 48.2
WR 27.5 27.4 32.2 28.4 31.8 35.6 41.7 37.8 40.5 53.2
Moreover, in Table 8, we compare our method and various baselines under the same setting on the AlpacaEval 2.0 benchmark, including SLiC-HF (Zhao et al., 2023), ORPO (Hong et al., 2024), CPO (Xu et al., 2024a), RRHF (Yuan et al., 2024), KTO (Ethayarajh et al., 2024), IPO (Azar et al., 2023), R-DPO (Park et al., 2024), and SimPO (Meng et al., 2024), where the results are from Meng et al. (2024), as well as the RPO (Adler et al., 2024) baseline that we implement. Our method outperforms the above algorithms by a considerable margin.
@@
In previous parts, our
@@
method is implemented by conditioning on the scalar reward values given by the judge models, either LLMs or reward models. We find that our approach is generalizable to settings of multi-dimensional rewards that correspond to different attributes, such as helpfulness and truthfulness. Specifically, we follow the setting from last part to construct the preference dataset by applying the ArmoRM reward model on the on-policy responses generated by Llama-3-8B-Instruct. Since ArmoRM is a multi-objective model that not only gives a scalar reward value but also predicts human-interpretable fine-grained attributes, we first select $5$ attributes (namely complexity, instruction following, honesty, helpfulness, and intelligence depth) that have the highest average coefficients on the UltraFeedback data. Then we relabel the data by conditioning on the $5$-dim reward and follow the implementation of using ArmoRM described in the last part. The resulting model achieves state-of-the-art within the Llama-3-8B-Instruct model family, surpassing the strong baselines including SimPO (Meng et al., 2024) that is trained also on on-policy data ranked by ArmoRM, and OpenChat (Wang et al., 2023a) fine-tuned with Conditioned-RLFT from the same Llama-3-8B-Instruct model.
@@
We further compared with additional conditional post-training baselines on the offline UltraFeedback dataset (i.e., without on-policy data), including DPA (Wang et al., 2024a), SteerLM (Dong et al., 2023b), and (Info)NCA (Chen et al., 2024). Since both baselines aim to optimize a user-controllable attribute-conditioned LLM that is optimal under diverse preference profiles with different coefficients of the reward’s attributes, in Figure 4, we plot the win rate curves of these methods under varying preference profiles, such as adjusting verbosity preferences as considered in Wang et al. (2024a). It can be observed that fine-tuned from Zephyr-SFT, our method achieves the best AlpacaEval 2.0 win rate. In addition to the comparison with the implemented RPO (Adler et al., 2024) in Table 8, we also report the performance of RPO fine-tuned on additional models including Qwen2-7B-Instruct and Gemma2-9B-It. As shown in Table 10, the implemented RPO is outperformed by our method across these models.
Figure 4: Comparison with DPA, SteerLM, and (Info)NCA.
LC Win Rate
Win Rate
Avg. Len.
Qwen+RPO 20.29 17.34 1704
Qwen+DPO (RA) 31.17 27.58 1789
Gemma+RPO 43.14 30.93 1413
Gemma+DPO (RA) 59.27 54.56 1872
@@ A.2 Theoretical Formulation
Goal-conditioned preference model. Consider a language model $\pi\in\Delta_{\mathcal{Y}}^{\mathcal{X}}$ that takes the prompt $x\in\mathcal{X}$ as input and outputs the response $y\in\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ are spaces of prompts and responses, respectively.Given the prompt $x\in\mathcal{X}$, a discrete probability distribution $\pi(\cdot\mid x)\in\Delta_{\mathcal{Y}}$ is generated, where $\Delta_{\mathcal{Y}}$ is the set of discrete distributions over $\mathcal{Y}$. We define the goal-conditioned reward function class as $\mathcal{R}\subset\{R(x,y,g):\mathcal{X}\times\mathcal{Y}\times\mathcal{G}\mapsto\mathbb{R}\}$, where $\mathcal{G}$ is the goal space. The goal-conditioned Bradley-Terry model (Bradley and Terry, 1952) for annotations is described as
@@
Performance metric. For the notational simplicity in the theoretical analysis, we denote by $R^{\star}$ the ground-truth goal-conditioned reward function. The alignment target is to maximize the expected true reward $R^{\star}$ conditioned on the optimal goal $g^{\star}\in\mathcal{G}$. Thus, we define the value function of any policy $\pi$ as
@@
Here we allow the prompt distribution $d_{0}(\cdot)$ to be different from that of the offline dataset distribution $\mu_{\bar{\mathcal{D}}}(\cdot)$, but is assumed to be known. In the meanwhile, we consider the policies that share the same support as the reference policy $\pi_{\mathrm{ref}}$ (Xiong et al., 2023), that is, we take a policy class $\Pi$ as
@@
Theoretical version of the reward-augmented DPO. We formulate the theoretical version of the reward-augmented DPO in Algorithm 8, where we add a SFT regularizer on the empirical objective to handle the issue distribution shift and analyze the bound on the suboptimality (Liu et al., 2024; Cen et al., 2024). One simple choice to define SFT policy $\pi_{\text{SFT}}$ is to utilize the chosen labels in the original preference dataset $\mathcal{D}$, that is,
@@ B.1 Setup
For hyperparameters, we utilize a KL regularization coefficient of $\beta=0.01$ in DPO, and we adopt the AdamW optimizer (Loshchilov, 2017). The batch size is set to $128$, with a learning rate of $75e-$ 7 and a warmup ratio of $0.1$. Furthermore, we observe that for models such as Qwen2-7B-Instruct and Gemma-2-9B-It on UltraFeedback, as well as Llama-3-8B-Instruct on on-policy data, both DPO and our proposed method yield improved performance when employing the conservative DPO (cDPO) technique (Mitchell, 2023). Consequently, for these models, we set the label smoothing hyperparameter from the Alignment Handbook (alignment_handbook2023) to $0.3$, while keeping it at 00 for the remaining models.
@@ B.2 Full Results
We also provide the full comparison results with reward-augmented methods in Table 12.
@@ B.3 More Ablations
Controllable Generation with Prompt.
$g=10$
$g=8$
$g=6$
UF ($g=10$)
LC WR 31.17 28.66 25.56 24.44
WR 27.58 25.57 18.88 20.75
In Table 13, we ablate how generations differ when changing the goal rewards in the system prompt. We observe that the AlpacaEval 2.0 scores of the Qwen2-7B-It+DPO (RA) model change accordingly as $g$ varies. However, using the same $g=10$ prompt during inference for the Qwen2-7B-It+DPO (UF) model fails to give competitive results, indicating that our method is superior not only because of the additional system prompt.

### v5

@@ Results.
Besides, we also evaluate the models on six academic multi-choice question-answering benchmarks, including GSM8K (Cobbe et al., 2021), GPQA (Rein et al., 2023), MUSR (Sprague et al., 2023), TruthfulQA (Lin et al., 2021), BBH (Suzgun et al., 2022), and ARC Challenge (Clark et al., 2018). To better reflect the capabilities of LLMs, we adopt various settings for these benchmarks, including zero-shot, few-shot, and few-shot Chain-of-Thought (CoT). The results are shown in Table 3.
@@ 6.2 Ablation Studies
We provide two pieces of evidence that our method can get more juice out of the preference data compared to directly applying DPO. Firstly, we evaluate SPPO (Wu et al., 2024) fine-tuned with DPO on UltraFeedback (UF). The results are shown in Table 4. Since the SPPO model is already trained on UltraFeedback from Gemma-2-9B-It, an additional round of DPO training with the same data significantly degrades its performance. In contrast, performing DPO on Reward-Augmented (RA) UltraFeedback results in substantial performance gains for SPPO, indicating that our method enhances the utility of the preference data.
@@
The second evidence is that after DPO, the implicit reward can be used to relabel and augment the same preference data. Specifically, after training Qwen2-7B-Instruct with DPO on UltraFeedback, we leverage the resulting model $\pi_{\text{DPO}}$ to calculate the implicit reward for each prompt $x$ and response $y$, i.e., $\widehat{r}=\beta(\log\pi_{\text{DPO}}(y\mid x)-\log\pi_{\text{Qwen}}(y\mid x))$. Then we perform DPO on Qwen2-7B-Instruct using the Implicit-Reward-Augmented (IRA) UltraFeedback. The results are shown in Table 5. We observe that augmenting the data with the implicit reward from the DPO (UF) model leads to superior performance even compared to augmenting the data with reward scores from the LLM judge, i.e., DPO (RA). This result highlights that DPO does not fully exploit the potential of the data. Moreover, this ablation demonstrates that our method is compatible with binarized preference datasets that only contain chosen and rejected response pairs, bypassing the need of judge models.
@@
In this part, we show that the success of our method is not merely due to the increased size of the training dataset. To illustrate this, we perform DPO on the
@@
dataset where reward augmentation is applied to the first half of the UltraFeedback data, which we denote as DPO (Half RA). By doing so, the reward-augmented data is of the same size as the original dataset, but with only half of the prompts and the corresponding responses being utilized. It can be observed from Table 6 that DPO (Half RA) outperforms fine-tuning on the whole UltraFeedback (UF) by a large margin and achieves comparable performance to applying reward augmentation across the entire UF dataset, which is denoted as DPO (RA).
Reward-Augmented Data Mitigates the Unlearning Issue.
We first demonstrate that DPO suffers from the limitation of unnecessarily unlearning high-quality rejected responses, as discussed in Section 3.1. Specifically, on the test set of UltraFeedback, we calculate the log probability of each rejected response for the Qwen2-7B-Instruct model, its DPO (UF) model, and our method DPO (RA). In Fig. 3, we plot the expected log probability for rejected responses
@@
whose reward scores $\geq 5$. We find that DPO substantially decreases the log probability of these high-quality rejected responses, confirming that the unlearning issue arises in practice. In contrast, our method alleviates this issue, although the probability is still slightly lower than the base model, which is proven to be DPO’s feature (Rafailov et al., 2024a; Zhang et al., 2024; Xu et al., 2024b).
@@
In previous parts, our method is implemented by conditioning on the scalar reward values given by the judge models, either LLMs or reward models. We find that our approach is generalizable to settings of multi-dimensional rewards that correspond to different attributes, such as helpfulness and truthfulness. Specifically, we follow the setting from last part to construct the preference dataset by applying the ArmoRM reward model on the on-policy responses generated by Llama-3-8B-Instruct. Since ArmoRM is a multi-objective model that not only gives a scalar reward value but also predicts human-interpretable fine-grained attributes, we first select $5$ attributes (namely complexity, instruction following, honesty, helpfulness, and intelligence depth) that have the highest average coefficients on the UltraFeedback data. Then we relabel the data by conditioning on the $5$-dim reward and follow the implementation of using ArmoRM described in the last part. The resulting model achieves state-of-the-art within the Llama-3-8B-Instruct model family, surpassing the strong baselines including SimPO (Meng et al., 2024) that is trained also on on-policy data ranked by ArmoRM, and OpenChat (Wang et al., 2023a) fine-tuned with Conditioned-RLFT from the same Llama-3-8B-Instruct model.
@@
We further compared with additional conditional post-training baselines on the offline UltraFeedback dataset (i.e., without on-policy data), including DPA (Wang et al., 2024a), SteerLM (Dong et al., 2023b), and (Info)NCA (Chen et al., 2024). Since both baselines aim to optimize a user-controllable attribute-conditioned LLM that is optimal under diverse preference profiles with different coefficients of the reward’s attributes, in Figure 4, we plot the win rates of these methods under varying preference profiles, such as adjusting verbosity preferences as considered in (Wang et al., 2024a). Fine-tuned from Zephyr-SFT, our method achieves the best AlpacaEval 2.0 win rate. We defer more ablations and baseline comparisons to Appendix B.3.
Figure 4: Comparisons with DPA, SteerLM, and (Info)NCA.
Controllable Generation with Prompt.
In Table 8, we ablate how generations differ when changing the goal rewards in the system prompt. We observe that the AlpacaEval scores of the Qwen2-7B-It+DPO (RA) model change accordingly as $g$ varies. However, using the same $g=10$ prompt during inference for Qwen2-7B-It+DPO (UF) fails to give competitive results, indicating that our method is superior not only because of the additional system prompt.
$g=10$
$g=8$
$g=6$
UF ($g=10$)
LC WR 31.17 28.66 25.56 24.44
WR 27.58 25.57 18.88 20.75
We refer the readers to Appendix B.3 for more ablations.
@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.2 Theoretical Formulation
Goal-conditioned preference model. Consider a language model $\pi\in\Delta_{\mathcal{Y}}^{\mathcal{X}}$ that takes the prompt $x\in\mathcal{X}$ as input and outputs the response $y\in\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ are spaces of prompts and responses, respectively. Given the prompt $x\in\mathcal{X}$, a discrete probability distribution $\pi(\cdot\mid x)\in\Delta_{\mathcal{Y}}$ is generated, where $\Delta_{\mathcal{Y}}$ is the set of discrete distributions over $\mathcal{Y}$. We define the goal-conditioned reward function class as $\mathcal{R}\subset\{R(x,y,g):\mathcal{X}\times\mathcal{Y}\times\mathcal{G}\mapsto\mathbb{R}\}$, where $\mathcal{G}$ is the goal space. The goal-conditioned Bradley-Terry model (Bradley & Terry, 1952) for annotations is described as
@@
Performance metric. For notational simplicity in the theoretical analysis, we denote by $R^{\star}$ the ground-truth goal-conditioned reward function. The alignment target is to maximize the expected true reward $R^{\star}$ conditioned on the optimal goal $g^{\star}\in\mathcal{G}$. Thus, we define the value function of any policy $\pi$ as
@@
Here we allow the prompt distribution $d_{0}(\cdot)$ to be different from that of the offline dataset distribution $\mu_{\bar{\mathcal{D}}}(\cdot)$, but it is assumed to be known. In the meanwhile, we consider the policies that share the same support as the reference policy $\pi_{\mathrm{ref}}$ (Xiong et al., 2023), that is, we take a policy class $\Pi$ as
@@
Theoretical version of the reward-augmented DPO. We formulate the theoretical version of the reward-augmented DPO in Algorithm 8, where we add an SFT regularizer on the empirical objective to handle the issue of distribution shift and analyze the bound on the suboptimality (Liu et al., 2024; Cen et al., 2024). One simple choice to define SFT policy $\pi_{\text{SFT}}$ is to utilize the chosen labels in the original preference dataset $\mathcal{D}$, that is,
@@ B.1 Setup
For hyperparameters, we tune the KL regularization coefficient $\beta$ within $[0.001,0.01,0.1]$ and batch size within $[64,128,256]$. We find that $\beta=0.01$ and a $256$ batch size yield the overall best performance for DPO across models. Our method uses the same hyperparameters as DPO. Besides, we adopt the AdamW optimizer (Loshchilov, 2017), with a learning rate of $75e-$ 7 and a warmup ratio of $0.1$. Furthermore, we observe that for models such as Qwen2-7B-Instruct and Gemma-2-9B-It on UltraFeedback, as well as Llama-3-8B-Instruct on on-policy data, both DPO and our proposed method yield improved performance when employing the conservative DPO (cDPO) technique (Mitchell, 2023). Consequently, for these models, we set the label smoothing hyperparameter from the Alignment Handbook (Tunstall et al., 2023a) to $0.3$, while keeping it at 00 for the remaining models.
@@ B.2 Full Results
We also provide the full comparison results with reward-augmented methods in Table 10, which corresponds to Figure 4 in the main text.
@@ B.3 More Ablations
Impact of the Accuracy of AI Feedback.
Llama-3- PairRM (0.4B) ArmoRM (8B)
8B-Instruct
DPO (UF)
DPO (RA+)
DPO (RA)
DPO (UF)
DPO (RA+)
LC WR 22.92 41.76 44.72 48.20 42.32 48.73
WR 23.15 45.79 44.70 53.17 42.79 45.36
More Comparisons with Baselines
In Table 12, we compare our method and various baselines under the same setting on the AlpacaEval 2.0 benchmark, including SLiC-HF (Zhao et al., 2023), ORPO (Hong et al., 2024), CPO (Xu et al., 2024a), RRHF (Yuan et al., 2024), KTO (Ethayarajh et al., 2024), IPO (Azar et al., 2023), R-DPO (Park et al., 2024), and SimPO (Meng et al., 2024), where the results are from (Meng et al., 2024), as well as the RPO (Adler et al., 2024) baseline that we implement. Our method outperforms the above algorithms by a considerable margin.
SLiC-HF ORPO CPO RRHF KTO IPO RPO R-DPO SimPO Ours
LC WR 26.9 28.5 28.9 31.3 33.1 35.6 40.8 41.1 44.7 48.2
WR 27.5 27.4 32.2 28.4 31.8 35.6 41.7 37.8 40.5 53.2
In addition to the comparison with RPO from (Adler et al., 2024) in Table 12, we also report the performance of RPO fine-tuned on additional models including Qwen2-7B-Instruct and Gemma2-9B-It. As shown in Table 13, the implemented RPO is outperformed by our method across these models.
LC Win Rate
Win Rate
Avg. Len.
Qwen+RPO 20.29 17.34 1704
Qwen+DPO (RA) 31.17 27.58 1789
Gemma+RPO 43.14 30.93 1413
Gemma+DPO (RA) 59.27 54.56 1872